Question 18 (3 points) A couple decides to purchase a beach condominium in Hilton Head Island, South Carolina. The agreed purchase price will be $290,000. The couple will pay 20% down on the condo and finance the remaining balance with a 30-year mortgage. The terms of the mortgage are 4.80% APR (with monthly compounding. What is the monthly payment for the mortgage? Format $1,234.45 as $1234.45 Your Answer: Answer Question 19 (3 points) Janice has $5,000 invested in a bank that pays 7.7% annually. funds to triple? How long will it take for her 1) 18.69 years 2)14.85 years 3) 11.33 years 4) 14.81 years 5) 13.03 years
Answer #1

Question 18

Down payment = 290000 * 20% = $58000

Remaining amount mortgaged = $290000 - $58000 = $232000

EMI = [P x R x (1+R)^N]/[(1+R)^N-1] = [232000*0.048/12*[(1+0.048/12)^30*12/(1+0.048/12)^30*12 - 1] = $1217.22

Question 19

FV = PV (1+r)^n

15000 = 5000(1+0.077)^n

Let us consider all options,

IF n = 18.69 years then 5000(1+0.077)^18.69 = $100645.7

IF n = 14.85 years then 5000(1+0.077)^14.85 = $ 15044.34

IF n = 11.33 years then 5000(1+0.077)^11.33 = $ 11587.08

IF n = 14.81 years then 5000(1+0.077)^14.81 = $ 14999.77 i.e. $15000

Therefore correct answer is 14.81 years